Calculate the resolving power for m/z = 283.980 if the width of half height is 0.076 m/z. Note: RP = m/m 1/2 or RP = m/Am
Answer #1

solution: resolving power = 3737

resolving power (RP) is a dimensionless ratio defined as the mobility drift time (td) divided by the width of the peak (Δtd): RP = td / Δtd

mobility drift time (td) = 283.980

width of the peak (Δtd) = 0.076

putting above value in the resolving power equation

RP = 283.980 / 0.076 = 3737